# CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models

### v3

@@ 3.1 TextUnlock
We provide an illustration of the TexUnlock approach in Figure 1 and a PyTorch-like pseudocode in Section A of the Appendix. It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section O in the Appendix for more information).
3.2 Zero-Shot Concept Bottleneck Models
Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed zero-shot CBMs. Note that at this stage, all model components (including the MLP) are frozen and no further training is performed. CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction. In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image. In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
Concept Discovery. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier. We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works (Rao2024DiscoverthenNameTC), and without loss of generality, we use the $Z=20K$ most common words in English (mostcommongoogle) as our concept set. These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset. We ablate on other concept sets in Appendix Section I for the interested reader. To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set. Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”). In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. Details of this procedure can be found in Section N of the Appendix. This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes. With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot. This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
Figure 2: Building zero-shot CBMs for any pretrained classifier. We first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in a zero-shot manner, which results in (c) our final CBM. Note that the concept bank only needs to be encoded once.
We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$. That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$. By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$. For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space. That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$. Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to. That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$. The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts. We provide an illustration in Figure 2(a).
Concept-to-class Prediction. The classifier $W^{con}$ takes in concept activations and outputs a distribution $S_{cn}$ over classes. We build $W^{con}$ in a zero-shot manner, where here zero-shot indicates that no training is required to map concept activations to classes. Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space. Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name. Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$. That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$. Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts. This process is shown in Figure 2(b). In total, the output distribution $S_{cn}$ of the CBM is defined as:
[…]
From Eq. 2 we make an interesting observation. Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other. Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs preserve the basic reasoning process of the original classifier. By this, we obtain zero-shot CBMs that discover concepts and build $W^{con}$ in a zero-shot manner for any classifier (Figure 2(c)). An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly. This makes our method highly flexible with respect to the chosen concept set.
4 Experiments
Model Top-1 Orig. $\Delta$ Model Top-1 Orig. $\Delta$
ResNet50 75.80 76.13 -0.33 ResNeXt50-32x4dv2 80.79 80.88 -0.09
[…]
ResNet101 77.19 77.37 -0.18 ViT-B/16 80.70 81.07 -0.37
[…]
DenseNet169 75.46 75.60 -0.14 Swin-Base 83.22 83.58 -0.36
EfficientNetv2-S 84.04 84.23 -0.19 Swinv2-Base 83.72 84.11 -0.39
EfficientNetv2-M 84.95 85.11 -0.16 BeiT-B/16 84.54 85.06 -0.52
[…]
ConvNeXt-Small 83.42 83.62 -0.20 DINOv2-B 84.40 84.22 +0.18
ConvNeXt-Base 83.88 84.06 -0.18 ConvNeXtV2-B 84.56 84.73 -0.17
[…]
Table 1: TextUnlock-ed visual classifiers maintain classification performance. Comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock. We use the ImageNet-1K benchmark dataset due to the widespread publicly available visual classifiers trained and evaluated on it. In Section F of the Appendix, we also report results on other datasets including Places365 (zhou2017places), EuroSAT (helber2019eurosat) and DTD (Cimpoi2013DescribingTI) showing that our method is also applicable to domain-specific and fine-grained datasets. We apply TextUnlock on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He2015DeepRL), Wide ResNets (Zagoruyko2016WRN), ResNeXts (Xie2016AggregatedRT), ShuffleNetv2 (Ma_2018_ECCV), EfficientNetv2 (Tan2021EfficientNetV2SM), Densely Connected Networks (DenseNets) (Huang2016DenselyCC), ConvNeXts (Liu2022ACF) and ConvNeXtv2 (Woo2023ConvNeXtVC). For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (dosovitskiy2021an), DINOv2 (oquab2024dinov), BeiT (bao2022beit), the hybrid Convolution-Vision Transformer CvT (Wu2021CvTIC), Swin Transformer (Liu2021SwinTH) and Swin Transformer v2 (Liu2021SwinTV). All models are pretrained on ImageNet-1K from the PyTorch (torchvision2016) and HuggingFace (wolf-etal-2020-transformers) libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe (PytorchNewTrainRecepie). Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder (wang2020minilm) as it is powerful, fast and efficient. We provide ablation studies on other text encoders in Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section C of the Appendix (see Section C.1 for the role and impact of the MLP and Section C.2 for its design hyperparameters). We also evaluate our transformed classifier robustness to prompt variations in Section H of the Appendix. For results on other models, we refer to Section L of the Appendix, and for implementation details, to Section J of the Appendix.
[…]
Effectivenss of Concept Interventions. Another common way to evaluate interpretability of CBMs is concept intervention. We report concept intervention results on our ZS-CBMs in Section E of the Appendix. In this way, we show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
Supervised CBMs
Method Model Top-1 Method Model Top-1
LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
Zero-Shot CBMs (Ours)
Method Model Top-1 Method Model Top-1
ZS-CBM ResNet50 73.9 ZS-CBM ViT-B/32 73.3
[…]
ZS-CBM WideResNet50 76.9 ZS-CBM Swinv2-Base 82.6
[…]
ZS-CBM DenseNet121 69.9 ZS-CBM DINOv2-B 82.6
[…]
ZS-CBM EfficientNetv2-M 83.9 ZS-CBM BeiT-L/16 86.2
[…]
Table 2: Our zero-shot CBMs outperform CLIP-based counterparts. Accuracy of Supervised and Zero-Shot CBMs (ZS-CBMs, ours) on ImageNet validation set. Similar backbones are color-coded. Best within backbone family is underlined, overall best is bold.
In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder. All examples use the LF-CBM concept set (oikarinen2023labelfree). By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab. Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature (Samek2019TowardsEA) regarding the “dumbbell” class. Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it. With our method, we can obtain a textual interpretation of the biases that the original classifier learns. We also provide qualitative examples of global class-wise concepts detected in Section P of the appendix.
Figure 3: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the on the x-axis.
5 Zero-Shot Image Captioning
We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models. Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space. Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier. We adapt the method introduced in ZeroCap (Tewel2021ZeroCapZI) for our purpose. Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features. We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li2021PrefixTuningOC) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual" tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. As this process is not the core contribution of our work, we leave more details to Section M of the appendix, which also includes a detailed illustration in Figure 2.
[…]
Figure 4: Qualitative examples of zero-shot image captioning.
Model B4 M R-L C S
ZeroCap 2.6 11.5 14.6 5.5
ConZIC 1.3 11.5 12.8 5.2
Ours
DenseNet161 1.50 10.2 20.4 15.8 6.3
ResNet50 1.43 10.2 20.3 15.9 6.2
WideResNet50 1.40 10.2 20.4 16.0 6.4
[…]
DINOv2-Base 1.50 10.7 21.0 17.3 6.7
[…]
BeiT-L/16 1.50 10.6 20.9 17.6 6.9
[…]
Table 3: Zero-Shot Image Captioning Performance
Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this hypothesis by applying compositional image captioning (Kulkarni2011BabyTU; Lu2018NeuralBT) with in-context learning. Specifically, a set of image-grounded verbs and concepts (e.g., attributes, objects) are first detected using the concept discovery step described in Section 3.2, and they are fed along with six in-context examples from the COCO captioning set to a language model, capable of imitating the style and structure of the in-context examples, in order to compose them into a natural sounding sentence. This also demonstrates the direct application of our approach to compositional image captioning. The last row of Table 3 with the superscript com shows that the (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M compared to baseline methods. We provide more details and results on other models on this paradigm of compositional image captioning in Section K of the Appendix. We further show how we can use this approach to generate captions tailored to any domain (see Section K.1 in the Appendix).
6 Conclusion
We introduced a method for transforming any frozen visual classification model into a Concept Bottleneck Model (CBM). We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart. This method further enables a CLIP-free, label-free conversion of the frozen visual classification model into a CBM, and further allows us to derive zero-shot concept bottleneck concept-to-class classifiers (without training the linear classifier), Applied on 40 different models, our method outperforms supervised CLIP-based CBMs and achieves state-of-the-art results. TextUnlock further allows us to perform zero-shot image captioning on any visual classification model, also achieving state-of-the-art results and surpassing zero-shot captioning baselines based on CLIP. Finally, as with any research work, this study has its own limitations which are discussed in Section B of the Appendix.
Appendix A Pseudocode of TextUnlock
We provide a PyTorch-like pseudocode of TextUnlock in Listing 1.
@@ Appendix A Pseudocode of TextUnlock
Appendix B Limitations

### v4

@@ 3.1 TextUnlock
We provide an illustration of the TexUnlock approach in Figure 1, and a PyTorch-like pseudocode of it in Listing LABEL:lst:pseudo. It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section 14 in the Appendix for more information).
@@
3.2 U-F2-CBMs
Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed U-F2-CBMs. Note that at this stage, all model components (including the MLP) are frozen and no further training is performed. A CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction. In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image. In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
Concept Discovery: We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which represent the classification weights of the newly formulated classifier. In CBMs, we are given a large set of textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works [44] and without loss of generality, we use the $Z=20K$ most common words in English [12] as our concept set. These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset. We ablate on other concept sets in Appendix Section 7 for the interested reader. To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set. Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”). In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. Details of this procedure can be found in Section 13 of the Appendix. This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes. With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot. This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
We use the same text encoder $T$ that generates the linear classifier $U$ to generate concept embeddings, by feeding each concept $z_{i}\in\mathcal{Z}$ to the text encoder $T$ to generate a concept embedding $c_{i}$. That is, $c_{i}=T(z_{i})$, $\forall i=1,\dots,Z$. By performing this for all $Z$ concepts, we obtain a concept embedding matrix $C\in\mathbb{R}^{Z\times m}$. For an image $I$, we extract its visual features $f$ and use the MLP to map them to $\tilde{f}$ which now lies in the text embedding space. That is, $\tilde{f}=\text{MLP}(f)$, and $\tilde{f}\in\mathbb{R}^{m}$. Since $C$ and the mapped visual features $\tilde{f}$ are now in the same space, we can query $\tilde{f}$ to find which concepts it responds to. That is, we perform concept discovery using the cosine similarity between $\tilde{f}$ and each row-vector in $C$. The concept activations are obtained by $\tilde{f}.C\in\mathbb{R}^{Z}$ and represent the activation score for each of the $Z$ concepts. Concepts that the model identifies in an image will produce a high activation score. We provide an illustration in Figure 2(a).
Model Top-1 Orig. $\Delta$
[…]
Table 1: TextUnlock-ed visual classifiers for CNNs and Transformers maintain classification performance. Comparison of our re-formulated classifiers for several models (remaining models in Appendix). Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
Concept-to-Class Prediction: The classifier $W^{con}$ takes the concept activations produced from the concept discovery stage, and outputs a distribution $S_{cn}$ over classes. We build $W^{con}$ in an unsupervised manner; here unsupervised indicates that no training is required to map concept activations to classes. Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space. Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name. Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$. That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$. Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts. This process is shown in Figure 2(b). While the operation is performed entirely in the text feature space, the text encoder serves as the classifier weight generator. Because the MLP learns to map image representations to the text-encoder space, it is effectively mapping images into the classifier weight space. In total, the output distribution $S_{cn}$ of the CBM is obtained by feeding the identified concept activations in the image obtained from the concept discovery stage, to $W^{con}$. That is,
[…]
From Eq. 2 we make an interesting observation. Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other. Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs preserve the basic reasoning process of the original classifier. By this, we obtain CBMs that discover concepts and build $W^{con}$ in an unsupervised manner for any classifier (Figure 2(c)). An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly. This makes our method highly flexible to the concept set.
Supervised CBMs
Method Model Top-1 Method Model Top-1
LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
Unsupervised, Label-Free, CLIP-Free CBMs (U-F2-CBM)
Method Model Top-1 Method Model Top-1
U-F2-CBM ResNet50 73.9 U-F2-CBM ViT-B/32 73.3
[…]
U-F2-CBM WideResNet50 76.9 U-F2-CBM Swinv2-Base 82.6
[…]
U-F2-CBM DenseNet121 69.9 U-F2-CBM DINOv2-B 82.6
[…]
U-F2-CBM EfficientNetv2-M 83.9 U-F2-CBM BeiT-L/16 86.2
[…]
Table 2: Our U-F2-CBMs outperform CLIP-based counterparts. Accuracy of Supervised and U-F2-CBMs (ours) on ImageNet validation set. Similar backbones are color-coded. Best within backbone family is underlined, overall best is bold.
4 Experiments
We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock. We use the most challenging CBM benchmark of ImageNet-1K dataset. There is also widespread publicly available visual classifiers trained and evaluated on ImageNet. In Section 5 of the Appendix, we also report results on other datasets including Places365 [64], EuroSAT [14] and DTD [6] showing that our method is also applicable to domain-specific, fine-grained datasets as well as datasets with a small number of classes. We apply TextUnlock on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [13], Wide ResNets [61], ResNeXts [58], ShuffleNetv2 [31], EfficientNetv2 [50], Densely Connected Networks (DenseNets) [17], ConvNeXts [28] and ConvNeXtv2 [56]. For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [10], DINOv2 [38], BeiT [4], the hybrid Convolution-Vision Transformer CvT [57], Swin Transformer [27] and Swin Transformer v2 [26]. All models are pretrained on ImageNet-1K from the PyTorch [32] and HuggingFace [55] libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe [53]. BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder [54] as it is fast and efficient. We provide ablation studies on other text encoders in Section 2 of the Appendix. Results for 17 classifiers are presented in Table 1, with the remaining 23 classifiers reported in Section 11 of the Appendix. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section 3 of the Appendix (see Section 3.1 for the role and impact of the MLP and Section 3.2 for its design hyperparameters). We also evaluate our transformed classifier robustness to prompt variations in Section 6 of the Appendix. For implementation details, we refer to Section 8 of the Appendix.
[…]
Other Datasets:
We experiment with other datasets and show that our method is applicable even to domain-specific, fine-grained datasets as well as datasets with a small number of classes. We report CBM results on Places365 (domain-specific to scenes, 365 classes), DTD (domain-specific to texture and fine-grained, 47 classes), and EuroSAT (domain-specific to satellite images, 10 classes). When a baseline method is also reported on that dataset, we include it. Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline. All baselines use the same concept set. Note also that all baselines train a supervised linear classifier, while our method derives the classifier in an unsupervised manner. Results are shown in Table 3. For Places 365, we can see that U-F2-CBM for DenseNet161 classifier (using the transformed ImageNet-only trained classifier with TextUnlock) outperforms supervised CLIP-based ResNet and ViTs CBM methods. The same applies for EuroSAT and DTD with our U-F2-CBM. Therefore, the experiments show that our method scales to those scenarios as well.
Dataset Method Model Acc (%)
Places365 DCLIP CLIP-ResNet50 37.90
DCLIP CLIP-ViT-B/16 40.30
LF-CBM CLIP-ResNet50 49.00
LF-CBM CLIP-ViT-B/16 50.60
CDM CLIP-ResNet50 52.70
CDM CLIP-ViT-B/16 52.60
Ours ResNet50 51.57
Ours DenseNet161 53.42
EuroSAT Baseline CLIP-ResNet50 86.27
Baseline CLIP-ViT-B/16 88.57
Ours ViT-B/16 93.65
Ours WideResNet101 94.12
Ours ResNet50 94.22
DTD Baseline CLIP-ResNet50 57.77
Baseline CLIP-ViT-B/16 61.86
Ours WideResNet101 66.97
Ours ViT-B/16 68.46
Ours ResNet50 68.88
Table 3: CBM results on Places365, EuroSAT, and DTD datasets.
Figure 3: Qualitative examples of our U-F2-CBM. We show the top-detected concepts, each with their corresponding importance score.
Concept Interventions:
Another common way to evaluate interpretability of CBMs is concept intervention. This follows metrics from the explainability literature [41]. In this way, we can show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions. As these metrics serve as complementary evaluations, we refer readers to Section 4 of the Appendix for the results.
In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder. All examples use the LF-CBM concept set [36]. By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab. Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [47] regarding the “dumbbell” class. Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it. With our method, we can obtain a textual interpretation of the biases that the original classifier learns. We also provide qualitative examples of global class-wise concepts detected in Section 15 of the appendix.
5 Zero-Shot Image Captioning
We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models. Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space. Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier. We adapt the method introduced in ZeroCap [51] for our purpose. Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features. We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-shelf pretrained language decoder model, denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [22] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. More details about ZeroCap is in Section 12 of the appendix.
[…]
Model B4 M R-L C S
ZeroCap 2.6 11.5 14.6 5.5
ConZIC 1.3 11.5 12.8 5.2
Ours
ResNet50 1.43 10.2 20.3 15.9 6.2
[…]
DINOv2-Base 1.50 10.7 21.0 17.3 6.7
BeiT-L/16 1.50 10.6 20.9 17.6 6.9
[…]
Table 4: Zero-Shot Image Captioning Performance
6 Conclusion
We introduced a method for transforming any frozen visual classification model into a CBM. We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart. This allows us to then produce an unsupervised, label-free and CLIP-free (U-F2) CBM, which outperforms supervised CLIP-based CBMs across 40 models. We also showed zero-shot image captioning as an additional application. Finally, as with any research work, this study has its own limitations, discussed in Section 1 of the Appendix.
Supplementary Material
1 Limitations